Hankins Corporation has 7.3 million shares of common stock outstanding, 265,000 shares of 4.5 percent preferred stock outstanding, par value of $100; and 150,000 bonds with a semiannual coupon rate of 5.4 percent outstanding, par value $2,000 each. The common stock currently sells for $64 per share and has a beta of 1.20, the preferred stock has a par value of $100 and currently sells for $92 per share, and the bonds have 16 years to maturity and sell for 106 percent of par. The market risk premium is 7.4 percent, T-bills are yielding 3.2 percent, and the company’s tax rate is 22 percent. a. What is the firm’s market value capital structure? (Do not round intermediate calculations and round your answers to 4 decimal places, e.g., .1616.) b. If the company is evaluating a new investment project that has the same risk as the firm’s typical project, what rate should the firm use to discount the project’s cash flows?
Answer #1
MV of equity=Price of equity*number of shares outstanding
MV of equity=64*7300000
=467200000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=2000*150000*1.06
=318000000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=92*265000
=24380000
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=467200000+318000000+24380000
=809580000
a. Weight of equity = MV of Equity/MV of firm
Weight of equity = 467200000/809580000
W(E)=0.5771
a. Weight of debt = MV of Bond/MV of firm
Weight of debt = 318000000/809580000
W(D)=0.3928
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 24380000/809580000
W(PE)=0.0301
Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 3.2 + 1.2 * (7.4)
Cost of equity% = 12.08
Cost of debt
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =16x2
2120 =∑ [(5.4*2000/200)/(1 + YTM/200)^k]     +   2000/(1 + YTM/200)^16x2
                   k=1
YTM = 4.856314936
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 4.856314936*(1-0.22)
= 3.78792565008
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 4.5/(92)*100
=4.89
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=3.79*0.3928+12.08*0.5771+4.89*0.0301
b.WACC =8.61%